Fermi National Laboratory


Accelerator Update for December 10, 2004 - December 13, 2004

Friday December 10
The day shift began with Operations monitoring store 3829, with the Antiproton Source (Pbar) stacking, with the Main Injector (MI) and Recycler conducting studies, and with MiniBooNE and Switchyard (SY) taking beam.

A MI expert began a slipstacking study at 11:05 AM.

SY experts established beam to the SY abort at 11:11 AM.

Tevatron (TeV) experts began an end-of-store study at 11:17 AM.

At 12:19 AM, operators reported that they were having trouble with the Pbar reverse proton tune up.

Operations terminated store 3829 at 1:37 PM. TeV experts began a study period.

SY experts established beam to the Meson target train at 1:42 PM.

SY experts established slow spill to Meson enclosure M01 at 4:18 PM.

Operations reported at 6:24 PM that the Recycler was losing beam. Later it was discovered to be a MECAR problem, which was corrected.

Operations put the TeV into shot setup at 10:50 PM, but at 11:33 PM both the TeV proton and antiproton kickers pulled the TeV abort.

Saturday December 11
The midnight shift began with the TeV in shot setup, with stacking on hold for the shot, and with MiniBooNE taking beam but having tuning problems.

At 1:18 AM, Operations reported having problems with the H- Source. Operators accessed the dome and discovered a tripped voltage breaker for the gas pulser. It reset okay and everything was back to normal by 1:30 AM.

Operations established store 3834 at 2:56 AM with an initial luminosity of 52.51E30. Pbar resumed stacking.

Operations reported at 3:41 AM that operators were sending beam to the SY dump.

Operations started a shot setup for antiprotons to Recycler at 5:07 AM. By 6:56 AM, the Recycler had taken 26mA. Pbar resumed stacking.

The Meson MCenter experimenters requested beam at 8:06 AM. Operations had beam to the experiment five minutes later.

Operators prepared MCenter for an access at 9:02 AM. Operators turned SY back on and began tuning beam to MTest at 2:47 PM. SY experts conducted studies until 6:39 PM.

Sunday December 12
The midnight shift began with Operations monitoring store 3834, with Pbar stacking, with Operations monitoring a Recycler stash, with SY experts conducting studies, and with MiniBooNE taking beam.

SY studies ended at 12:53 AM.

Operations terminated store 3834 at 7:03 AM.

Operations put the TeV through a dry squeeze at 7:29 AM.

Operations put the TeV into shot setup at 7:51 AM.

Operations had injected protons and antiprotons into the TeV and started the TeV squeeze when the store was lost at 10:38 AM when the TeV quenched at sector F. Experts investigated and found a separator had sparked while trying to change its polarity.

Operators put the TeV separator in condition mode at 5 PM.

Operators put the TeV through a dry squeeze at 5:54 PM.

At 6:28 PM, Operations reported that some of the separator polarities were wrong and so conditioning had to be started over.

At 7:36 PM, Operations reported that Pbar was losing antiprotons due to a bad setting sent to Accumulator RF station ARF1. With the help of experts, Pbar stacking started looking better by 9:56 PM.

Operations established a 14x0 study-store at 10:31 PM.

Monday December 13
The midnight shift began with Pbar stacking, with Operations monitoring a Recycler stash, with Operations conditioning TeV separators, and with MiniBooNE, MCenter, and MTest taking beam.

At 12:33 AM, operators discovered more TeV separator polarity problems. The separator at A49 seems to be switching its own polarity, but not when it's requested. Eventually, operators got a hold of an expert to come in and look at the problem.

MTest experimenters finished with beam at 2:39 AM.

At 5:50 AM, Operations experienced problems with the H- Source. Operators accessed and found a problem with a power supply. It reset okay.

The TeV expert looking at the separator problem reported at 6:56 AM that he replaced a bad local/remote card. Operators put the separator into conditioning mode.

Plans for Monday
The plan for today is to establish a store this morning. Pbar may be allowed to conduct reverse proton studies this evening if everything is okay with the store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/13/2004   email Fermilab